# Carbon Emission Flow Tracing: Fast Algorithm and California Grid Study

### v1

@@ 1 Introduction
Achieving low-carbon or net-zero emissions in power systems is a critical priority for tackling climate change and enabling a sustainable energy future. In recent years, many countries, states, and companies have set ambitious decarbonization goals for the next two to three decades (Rockström et al., 2017; Papadis and Tsatsaronis, 2020; Newsom, 2022; Sustainability, ). As sectors such as transportation, heating, and computing become increasingly electrified—with data centers consuming rapidly growing amounts of power—the demand for clean electricity continues to rise (Xie et al., 2024). Accurate carbon emission quantification from the power grid serves a wide range of important applications: supporting emissions accounting, informing operational assessments and green tariff design, and guiding investment decisions on the siting and sizing of carbon-free resources (Ling et al., 2024). Reflecting growing public and institutional interest, a few independent system operators have begun publishing system-level emission data, such as total greenhouse gas emissions from CAISO (CAISO, c) and marginal emission rates from PJM (PJM Inside Lines, 2021).
@@
Fine-grained, high-quality emission data are essential for decarbonizing the power grid. Geographically distributed generators vary widely in technologies and emissions, and the electricity they produce is transported through the interconnected transmission and distribution networks to end users. Consequently, electricity consumed at a given location can reflect a different mix of sources (kir, 2002) and emission rates (Cho et al., 2025; Gorka et al., 2025). This makes it critical to understand when and where emissions occur to inform operational decisions and decarbonization strategies. High-resolution emission data are increasingly needed by policymakers, grid operators, and electricity end users who aim to track and reduce carbon impacts. These stakeholders can leverage high-fidelity spatiotemporal emission data to evaluate climate policies (Li and Lin, 2013), provide the right incentives (Li et al., 2022), and design emission-aware planning and operational models (Lindberg et al., 2021; Shao and Yu, 2025) that support a just and efficient clean energy transition.
@@
However, analyzing fine-grained spatiotemporal emission patterns remains challenging due to the inherent complexity of the power grid. The network structure, time-varying load and generation profiles, and physical limits on line flows make it difficult to attribute emissions from generators to specific locations and times. As a result, most estimates are coarse, system-level aggregates based on total energy mix, with limited spatial resolution. This limitation has real consequences. Some emission reduction strategies can lead to unintended effects—for instance, (Gorka et al., 2024) shows that load reduction at certain nodes may actually increase overall system emissions. Similarly, large electricity consumers such as manufacturing and IT companies face challenges in quantifying Scope 2 emissions—indirect emissions from purchased electricity—due to the lack of time- and location-specific data tied to electricity usage (Wyatt, 2024).
@@
Two key metrics in this context are the locational average emission rate (LAE), which captures total emissions per MWh over the entire electricity usage, and the locational marginal emission rate (LME), which reflects the emissions impact of a small change in local demand. The former, LAE, is particularly useful for quantifying total emissions of end users, supporting Scope 2 emission reporting and sustainability goal tracking. The latter parallels the concept of locational marginal price (LMP) and captures the sensitivity of emissions to local demand changes, making it well-suited for demand-side response, load shifting, and carbon reduction incentive design.
@@
Figure 1: Proposed framework for carbon tracing. Our approach is fully compatible with ACOPF, and can efficiently calculate the nodal average and marginal emission rate, which is validated and scaling up for the whole CAISO system.
@@
In this work, we develop a novel carbon tracing algorithm that supports both AC and DC power flow models and efficiently computes LAE and LME. Our method takes the grid topology, nodal demand, and renewable generation as input, and computes the exact contribution of each generator to each bus and transmission line. One key challenge in carbon tracing arises from cycle flows in the directed power flow graph, which can result from modeling artifacts, ACOPF non-convexity, or topological errors. We address this challenge by identifying and eliminating cycles using strongly connected components (SCCs), collapsing each SCC into a supernode. This transformation yields a directed acyclic graph, enabling an order of magnitude faster computation of nodal emission rates with linear complexity. The overall procedure is illustrated in Fig. 1.
We validate our approach through a large-scale case study using the California Test System (CATS) model (Taylor et al., 2024). CATS provides geographically accurate, realistic-yet-public grid topology data and is compatible with CAISO generation and demand records. Using approximated nodal demand and renewable generation, we solve the optimal power flow (OPF) problem to estimate power dispatch decisions, line power flows, and compute nodal emission rates. To validate the fidelity of our approach, we aggregate our nodal estimates and compare them to CAISO’s reported system-wide emissions. Across a full year of hourly data, our method achieves total emission estimates within 6% of CAISO’s published values.
@@
To the best of our knowledge, this is the first study to conduct nodal emission analysis on a realistic ISO-level power system. Moreover, our results indicate that there could be over 100 times difference in terms of county-level emission rate, highlighting substantial spatial disparities driven by generation mix and power flow. The methods and findings presented here offer a quantitative foundation for future research on locational emission analysis, including renewable impact assessment, siting of low-carbon resources, and generation expansion planning. Together, these tools can inform actionable strategies for emission reduction and sustainable energy transitions.
@@ 1.1 Related Works
More recently, LAEs and LMEs have gained significant attention as an effective metric for enabling load shifting and emissions reduction in energy-intensive applications (Lindberg et al., 2021). On the one hand, model-based approaches explicitly calculate such metrics using dispatch models. For instance, to calculate LMEs, implicit differentiation is adopted to take dynamic constraints such as storage devices and ramp-constraint generators into consideration (Valenzuela et al., 2023). While emission sensitivity analysis for LMEs and LAEs based on market clearing model are also studied (Lu et al., 2024; Kang et al., 2015). Model-based approaches in the literature often require matrix inversion, which is both computationally expensive and applicable mainly for linearized power flow. On the other hand, data-driven LME estimation methods studied in (Mayes et al., 2024; Shao and Yu, 2025) can face generalization challenges across varying topology and load conditions.
While most prior efforts address the quantification of LMEs and LAEs separately, our approach provides a unified framework based on physical power flow tracing that enables the computation of both metrics. Last but not least, the methods reviewed above—including ours—fall under the category of location-based emission accounting. In contrast, market-based approaches (Hulshof et al., 2019; Arellano and Carrión, 2023) estimate carbon emissions based on contractual arrangements for electricity procurement, such as Renewable Energy Certificates (RECs) or Power Purchase Agreements (PPAs), which is beyond the scope of this work.
@@
Carbon Tracing and Carbon Accounting: Pioneering works have established country-wide, real-time emission accounting frameworks. However, many do not model the exact physical power and carbon flows (Corradi, 2021), or are limited to estimating emissions at the balancing-authority or broader regional levels. For example, US-wide and European grid emission tracking are explored in (de Chalendar et al., 2019) and (Tranberg et al., 2019), respectively, but these studies rely on macro-level flow models rather than detailed transmission-level flows. Other methods (Chen et al., 2024a; Kang et al., 2015; Bialek, 1996) apply proportional sharing principles to allocate generator output and use matrix inversion to estimate the share of each generator’s power—and its associated emissions—attributed to each load.
@@ 2.1 Overview of the Proposed Framework
In this work, we aim to quantify each node’s carbon emission rate from power usage. At a given time snapshot, our approach takes the load vector and solves a general optimal power flow (OPF) problem using standard formulations and solvers. Once a feasible solution for generation dispatch and line flows is obtained, our algorithm traces each generator’s contribution to each demand node. To achieve so, we implement a tree-based algorithm that recursively, until all generators are enumerated, so that power generator can be fully allocated and traced to each demand node.
The full workflow is illustrated in Figure 1, and we particularly evaluate our algorithm for the whole California grid based on CATS model (Taylor et al., 2024). We begin with grid topology, renewable generation and line parameters acquired from CAISO and CATS repositories. Nodal load and renewables data are then fed into the OPF solver to obtain corresponding generation and line power flow values (Section 2.2). Once OPF instances are solved with feasibility and optimality guarantee, our carbon tracing algorithm computes the exact nodal power supply contributed by each generator with cycle flows considered (Sections 3.1 and 3.2). Locational emission rates are then derived based on EPA (Environmental Protection Agency, 2024) and EIA (US Energy Information Administration, 2024) actual generator’s emission rate datasets (see Table 3 in Appendix).
@@ 2.2 OPF and Power Dispatch
Once OPF problem (1) is solved, the nodal generations $\mathbf{p}^{g}$ and line flow $\mathbf{p}_{l}$ are revealed to the system operators. We denote the directed graph given by the OPF solution as $G^{\text{OPF}}$. Also use $\gamma_{k}$ denote the carbon emissions rate of generator $k$, then the total emissions associated with this generator can be quantified as $\gamma_{k}\cdot p_{k}^{g}$. We are interested in quantifying how much emissions are transported from generation side to electricity consumption through the solved power flows $p_{ij}$ with emission rate $\gamma_{ij}$.
@@ 2.3 Flow-Based Locational Emission Accounting
Our goal is to find the carbon emissions associated with each load node. We want to compute the LAE rate $\delta(p_{i}^{d})$ and LME rate $\mu(p_{i}^{d})$. Both rates have a unit of lbs CO2/MWh. To achieve this goal, we find it possible to follow the physical interpretations of average carbon emission rate $\delta(p_{i}^{d})$ via taking the division of nodal-level net emission by power demand $p_{i}^{d}$. The key for computing the total carbon emission is to find generator $k$’s contribution $p_{i}^{d}(p_{k}^{g})$ to supply $p_{i}^{d}$ in MW [^3] Throughout the paper, we use $p_{i}(x)$ to denote the power contributed by $x$ to node $i$, where $x$ denotes either generator or inflow. Similar definitions hold for line flow $f_{l}(x)$, then we can compute the total emission associated with $p_{i}^{d}$ as $e(p_{i}^{d})=\sum_{k=1}^{K}\gamma_{k}\cdot p_{i}^{d}(p_{k}^{g})$. Thus, each node’s average carbon emission rate (with $p_{i}^{d}>0$) is computed as,
@@
Figure 3: A 2-Generator, 1-Load illustration of average and marginal emission rate. When the load is small (left), cheaper renewables from Bus 2 are dispatched with 0 emissions. When the renewable generation is binding at high load times, both Bus 3’s average and marginal emission rate are changed due to traditional generator’s output.
In Fig. 3, we show a 3-bus toy example to illustrate how locational emission rates are quantified and varied due to load and generation conditions. In this simplified case, we assume the renewable generation are much cheaper but with fluctuating generations. Once power dispatch is determined based on (1), the proposed method will firstly determine each generator’s power contribution to each load and then trace generator’s contribution to nodal and line flow emissions.
@@ 3.1 Carbon Tracing Algorithm
Fig. 4 provides an illustrative example by solving the CAISO network’s full ACOPF using CATS model. Snippet of the whole network is shown for visualization purpose. In Fig. 4(b), $n1$ is connected to the generator $G2$ and has no other inflow power lines (noting the direction of the solved power flow, $n1$ only has outflows but not inflows), thus $\text{GNDF}_{G2}^{n1}$ = 1, meaning bus $n1$ is $100\%$ supplied by $G2$. $n1$ transfers power to bus $n3$ and bus $n4$. Based on Equ (5), line $f_{n1,n4}$ and $f_{n1,n3}$ are also $100\%$ supplied by $G2$, leading to $\text{GLDF}_{G2}^{n1,n3}=\text{GLDF}_{G2}^{n1,n4}=\text{GNDF}_{G2}^{n1}=1$. While at bus $n4$, power mix happens, requiring to know the quantity and power mix of $f_{n1,n4}$ and $f_{n3,n4}$ to calculate $\text{GNDF}_{G2}^{n4}=\frac{\text{GLDF}_{2}^{n1,n4}\cdot f_{n1,n4}+\text{GLDF}_{2}^{n3,n4}\cdot f_{n3,n4}}{f_{n1,n4}+f_{n3,n4}}$. With similar logic, we can analyze the rest of the buses in the topological order by iteratively updating these two matrices.
@@
To initialize this iterative process, at the initialization stage, we need to identify the load buses reachable by each generator. To achieve so, we treat each generator $k,k=1,...,K$ as the root node, and use Depth First Search (DFS) to get the buses and lines reachable from this $k$-th generator, which creates a mapping $\mathcal{M}(j)$ from bus $j$ to the generator set. In the mapping, the key is the bus index $j$, and each value is a set of generators supporting this bus’s power demand. In the same DFS recursion, we can also get the backtracking map. To initialize the GNDF and GLDF matrices and reduce complexity in subsequent calculations for each matrix entry, we flag $\text{GNDF}_{k}^{j}$ to 1 if bus $j$ is directly reachable by generator $k$ in the directed graph, and bus $j$ has no other inflow lines, e.g., $\mathcal{L}_{in}^{j}=\emptyset$. Similarly, $\text{GLDF}_{k}^{l}=1$ if line $l$ is flowing out of generator $k$, and $\mathcal{L}_{in}^{k}=\emptyset$. To initiate the iterative process, we need to store buses with zero in-degrees in an initial queue $\mathcal{Q}$. By doing so, in the latter stages we pop buses out of the queue in topological order and process them accordingly.
[…]
@@
4:  $\bm{\gamma}\leftarrow$ Generation emission rates vector
@@
6:  $\delta_{j}\leftarrow$ node $j$’s average emission rate $\forall j$
@@
[…]

### v2

@@ 1 Introduction
Achieving low-carbon or net-zero emissions in power systems is a critical priority for tackling climate change and enabling a sustainable energy future. In recent years, many countries, states, and companies have set ambitious decarbonization goals for the next two to three decades (Rockström et al., 2017; Papadis and Tsatsaronis, 2020; Newsom, 2022; Sustainability, ). As sectors such as transportation, heating, and computing become increasingly electrified—with data centers consuming rapidly growing amounts of power—the demand for clean electricity continues to rise (Xie et al., 2024). Accurate carbon emission quantification serves a wide range of important applications: supporting emissions accounting, informing operational assessments and green tariff design, and guiding investment decisions on the siting and sizing of carbon-free resources (Ling et al., 2024). Reflecting growing public and institutional interest, a few independent system operators have begun publishing system-level emission data, such as total greenhouse gas emissions from CAISO (CAISO, c) and marginal emission rates from PJM (PJM Inside Lines, 2021).
@@
Fine-grained, high-quality emission data are critical for decarbonizing the power grid by guiding generation dispatch, load shifting and emission-oriented planning (Cheng et al., 2022). Recent greenhouse gas (GHG) Protocol’s proposed revisions to Scope 2 guidance also highlights the requirements of such emission tracking capabilities under new carbon accounting standards (Protocol, ). Geographically distributed generators vary widely in technologies and emissions, and the electricity they produce is transported through the interconnected transmission and distribution networks to end users. Consequently, electricity consumed at a given location can reflect a different mix of sources (Kirschen et al., 1997) and emission rates (Cho et al., 2025; Gorka et al., 2025). This makes it critical to understand when and where emissions occur to inform operational decisions and decarbonization strategies. More recently, large electricity consumers such as manufacturing and IT companies face challenges in quantifying Scope 2 emissions—indirect emissions from purchased electricity—due to the lack of time- and location-specific data tied to electricity usage (Wyatt, 2024). High-resolution emission data are increasingly needed by policymakers, grid operators, and electricity end users who aim to track and reduce carbon impacts. These stakeholders can leverage high-fidelity spatiotemporal emission data to evaluate climate policies (Li and Lin, 2013), provide the right incentives (Li et al., 2022), and design emission-aware planning and operational models (Lindberg et al., 2021; Shao and Yu, 2025) that support a just and efficient clean energy transition.
@@
However, analyzing fine-grained spatiotemporal emission patterns remains challenging due to the inherent complexity of the power grid. The network structure, time-varying load and generation profiles, and physical limits on line flows make it difficult to attribute emissions from generators to specific locations and times. As a result, most estimates are coarse, system-level aggregates based on total energy mix, with limited spatial resolution. This limitation has real consequences. Some emission reduction strategies can lead to unintended effects—for instance, (Gorka et al., 2024) shows that load reduction at certain nodes may actually increase overall system emissions.
@@
Two key metrics in this context are the locational average emission rate (LAE) capturing total emissions per MWh over the entire electricity usage, and the locational marginal emission rate (LME) reflecting the emissions impact of a small change in local demand. The former, LAE, is particularly useful for quantifying total emissions of end users, supporting Scope 2 emission reporting and sustainability goal tracking (Protocol, ). The latter parallels the concept of locational marginal price (LMP) and captures the sensitivity of emissions to local demand changes, making it well-suited for demand-side response, load shifting, and carbon reduction incentives.
@@
Figure 1: Proposed framework for carbon tracing, fully compatible with ACOPF and applicable for nodal average and marginal emission rate.
@@
In this work, we develop a novel carbon tracing algorithm that supports both AC and DC power flow models and efficiently computes LAE and LME. Our method takes the grid topology, nodal demand, and renewable and conventional generation as input, and computes the exact contribution of each generator to each bus and transmission line. One key challenge in carbon tracing arises from cycle flows in the directed power flow graph, which can result from modeling artifacts, ACOPF non-convexity, or topological errors. We address this challenge by identifying and eliminating cycles using strongly connected components (SCCs), collapsing each SCC into a supernode. This transformation yields a directed acyclic graph, enabling an order of magnitude faster computation of nodal emission rates with linear complexity. We validate our approach through a large-scale case study using the California Test System (CATS) model (Taylor et al., 2024). CATS provides geographically accurate, realistic-yet-public grid topology data and is compatible with CAISO generation and demand records. Using approximated nodal demand and renewable generation, we solve the optimal power flow (OPF) problem to estimate power dispatch decisions, line power flows, and compute nodal emission rates. To validate the fidelity of our approach, we aggregate our nodal estimates and compare them to CAISO’s reported system-wide emissions. Across a full year of hourly data, our method achieves total emission estimates within 6% of CAISO’s published values. Fig. 1 illustrates the overall procedure. The major contributions are summarized as follows:
  • We develop a novel and computationally efficient carbon tracing approach, which is built upon physical power flow, and can be applied to both locational average and marginal emission rate calculations.
@@
  • Proposed method is fully compatible with full AC power flow, which reduces SCCs and avoids matrix inversions. The method and findings presented here offer a quantitative foundation for practices on locational emission analysis, including renewable impact assessment, siting of low-carbon resources, and generation expansion planning. Together, these tools can inform actionable strategies for emission reduction and sustainable energy transitions.
  • To the best of our knowledge, we provide the first study to conduct nodal emission analysis on a realistic ISO-level power system, and can provide hourly-matched and highly accessible carbon tracing. Results indicate there could be over 100 times difference in terms of county-level emission rate, highlighting substantial spatial disparities driven by generation mix and power flow.
@@ 1.1 Related Works
More recently, LAEs and LMEs have gained significant attention as effective metrics for enabling load shifting and emission reduction in energy-intensive applications (Lindberg et al., 2021). On the one hand, model-based approaches explicitly calculate such metrics using dispatch models. For instance, to calculate LMEs, implicit differentiation take dynamic constraints like storage and ramp-constraint generators into consideration (Valenzuela et al., 2023). While emission sensitivity analysis for LMEs and LAEs based on market clearing model are also studied (Lu et al., 2024; Kang et al., 2015). Model-based approaches in the literature often require matrix inversion, which is both computationally expensive and applicable mainly for linearized power flow. On the other hand, data-driven LME estimation methods can face generalization challenges across varying topology and load (Mayes et al., 2024; Shao and Yu, 2025).
While most prior efforts address the quantification of LMEs and LAEs separately, our approach provides a unified framework based on physical power flow tracing. Last but not least, methods reviewed above—including ours—fall under the category of location-based emission accounting. In contrast, market-based approaches (Hulshof et al., 2019) estimate carbon emissions based on contractual arrangements for electricity procurement, such as Renewable Energy Certificates (RECs) or Power Purchase Agreements (PPAs), which is beyond the scope of this work.
@@
Carbon Tracing and Carbon Accounting: Pioneering works have established country-wide, real-time emission accounting frameworks. However, many do not model the exact physical power and carbon flows (Corradi, 2021), or are limited to estimating emissions at the balancing-authority or broader regional levels. For example, US-wide and European grid emission tracking are explored in (de Chalendar et al., 2019) and (Tranberg et al., 2019), respectively, but these studies rely on macro-level flow models rather than detailed transmission-level flows. Other methods (Chen et al., 2024a; Kang et al., 2015) apply proportional sharing principles to allocate generator output and use matrix inversion to estimate the share of each generator’s power—and its associated emissions—attributed to each load.
@@ 2.1 Overview of the Proposed Framework
In this work, we aim to quantify each node’s carbon emission rate from power usage. At a given time snapshot, our approach takes the load vector and solves a general optimal power flow (OPF) problem using standard formulations and solvers. Once a feasible solution for generation dispatch and line flows is obtained, our algorithm traces each generator’s contribution to each demand node. To achieve so, we implement a tree-based algorithm recursively, until all generators are enumerated. This allows the power from every generator to be allocated and attributed to each demand node.
The full workflow is illustrated in Figure 1, and we particularly evaluate our algorithm for the whole California grid based on the open-sourced CATS model (Taylor et al., 2024), which preserves transmission line accuracy and reflects grid fidelity at the same time. We begin with grid topology, renewable generation and line parameters acquired from CAISO and CATS repositories. Nodal load and renewables data are then fed into the OPF solver to obtain corresponding generation and line power flow values (Section 2.2). Once OPF instances are solved with feasibility and optimality guarantee, our carbon tracing algorithm computes the exact nodal power supply contributed by each generator with cycle flows considered (Sections 3.1 and 3.2). Locational emission rates are then derived based on EPA (Environmental Protection Agency, 2024) and EIA (US Energy Information Administration, 2024) actual generator’s emission rate datasets (see Table 3 in Appendix).
@@ 2.2 OPF and Power Dispatch
Once OPF problem (1) is solved, the nodal generations $\mathbf{p}^{g}$ and line flow $\mathbf{p}_{l}$ are revealed to the system operators. We denote the directed graph given by the OPF solution as $G^{\text{OPF}}$. Let $\gamma_{k}$ denote the carbon emissions rate of generator $k$, then the total emissions associated with this generator can be quantified as $\gamma_{k}\cdot p_{k}^{g}$. We are interested in quantifying how much emissions are transported from generation to consumption through solved power flows $p_{ij}$ with emission rate $\gamma_{ij}$.
@@ 2.3 Flow-Based Locational Emission Accounting
Our goal is to find the carbon emissions associated with each load node. We want to compute the LAE rate $\delta(p_{i}^{d})$ and LME rate $\mu(p_{i}^{d})$. Both rates have a unit of metric tonnes CO2/MWh. To achieve this goal, we find it possible to follow the physical interpretations of average carbon emission rate $\delta(p_{i}^{d})$ via taking the division of nodal-level net emission by power demand $p_{i}^{d}$. The key for computing the total carbon emission is to find generator $k$’s contribution $p_{i}^{d}(p_{k}^{g})$ to supply $p_{i}^{d}$ in MW [^3] Throughout the paper, we use $p_{i}(x)$ to denote the power contributed by $x$ to node $i$, where $x$ denotes either generator or inflow. Similar definitions hold for line flow $f_{l}(x)$, then we can compute the total emission associated with $p_{i}^{d}$ as $e(p_{i}^{d})=\sum_{k=1}^{K}\gamma_{k}\cdot p_{i}^{d}(p_{k}^{g})$. Thus, each node’s average carbon emission rate (with $p_{i}^{d}>0$) is computed as,
@@
Figure 3: A 2-Generator, 1-Load illustration of average and marginal emission rate.
In Fig. 3, we show a 3-bus toy example to illustrate how locational emission rates are quantified and varied due to load and generation conditions. In this simplified case, we assume the renewable generation are much cheaper but with fluctuating generations. When the load is small (left), cheaper renewables from Bus 2 are dispatched with 0 emissions. When the renewable generation is binding at high load times, both Bus 3’s average and marginal emission rate are changed due to traditional generator’s output. Once power dispatch is determined based on (1), the proposed method will firstly determine each generator’s power contribution to each load and then trace generator’s contribution to nodal and line flow emissions.
@@ 3.1 Carbon Tracing Algorithm
Fig. 4 provides an illustrative example from the full ACOPF of the CAISO network based on the CATS model. Snippet of the whole network is shown for visualization purpose. In Fig. 4(b), $n1$ is connected to the generator $G2$ and has no other inflow power lines (noting the direction of the solved power flow, $n1$ only has outflows but not inflows), thus $\text{GNDF}_{G2}^{n1}$ = 1, meaning bus $n1$ is $100\%$ supplied by $G2$. $n1$ transfers power to bus $n3$ and bus $n4$. Based on Equ (5), line $f_{n1,n4}$ and $f_{n1,n3}$ are also $100\%$ supplied by $G2$, leading to $\text{GLDF}_{G2}^{n1,n3}=\text{GLDF}_{G2}^{n1,n4}=\text{GNDF}_{G2}^{n1}=1$. While at bus $n4$, power mix happens, requiring to know the quantity and power mix of $f_{n1,n4}$ and $f_{n3,n4}$ to calculate $\text{GNDF}_{G2}^{n4}=\frac{\text{GLDF}_{2}^{n1,n4}\cdot f_{n1,n4}+\text{GLDF}_{2}^{n3,n4}\cdot f_{n3,n4}}{f_{n1,n4}+f_{n3,n4}}$. With similar logic, we can analyze the rest of the buses in the topological order by iteratively updating these two matrices.
@@
To initialize this iterative process, at the initialization stage, we need to identify the load buses reachable by each generator. To achieve so, we treat each generator k,k=1,,Kk,k=1,...,K as the root node, and use Depth First Search (DFS) to get the buses and lines reachable from this $k$-th generator, which creates a mapping $\mathcal{M}(j)$ from bus $j$ to the generator set. In the mapping, the key is the bus index $j$, and each value is a set of generators supporting this bus’s power demand. In the same DFS recursion, we can also get the backtracking map. To initialize the GNDF and GLDF matrices and reduce complexity in subsequent calculations for each matrix entry, we flag $\text{GNDF}_{k}^{j}$ to 1 if bus $j$ is directly reachable by generator $k$ in the directed graph, and bus $j$ has no other inflow lines, e.g., $\mathcal{L}_{in}^{j}=\emptyset$. Similarly, $\text{GLDF}_{k}^{l}=1$ if line $l$ is flowing out of generator $k$, and $\mathcal{L}_{in}^{k}=\emptyset$. To initiate the iterative process, we need to store buses with zero in-degrees in an initial queue $\mathcal{Q}$. By doing so, in the latter stages we pop buses out of the queue in topological order and process them accordingly.
For each node $i\in\mathcal{N}$, we also compute and store the in-degrees $\text{In\_Deg}(i)\geq 0$ beforehand. In order to identify the starting point for calculating each entry in GNDF matrix, we utilize Kahn’s algorithm (Kahn, 1962) for topological sorting. It is a classic algorithm for deciding a valid ordering of tasks with dependencies on the directed acyclic graph, and is capable of processing very large networks. The process starts with nodes of $\text{In\_Deg}(\cdot)=0$ first, and remove the outflow lines of these nodes. The iterative process based on topological ordering is illustrated in Algorithm 1.
@@
3: Initialize: $\mathcal{Q}\leftarrow$ a queue of nodes with $\text{In\_Deg}(\cdot)=0$
@@
[…]
@@
The algorithm stops when there is no more bus in the queue $\mathcal{Q}$, i.e. there is no bus with $\text{In\_Deg}=0$ anymore. Kahn’s algorithm stops until it processes all the nodes in a directed acyclic graph (DAG). All the buses proportion values and line proportion values in our matrices are calculated as long as our graph topology is DAG. In the DCOPF case, previous works show that no cycle flow exist (Chen et al., 2024b, 2022), making our algorithm directly applicable. While in the solutions given by ACOPF, there may exist directed cycle flows, where there exists a non-empty directed trail in which the first and last nodes are equal. To make our carbon tracing algorithms apply to this general setting, we design a cycle detection and elimination framework in the next subsection.